# Parameters vs FLOPs: Scaling Laws for Optimal Sparsity for Mixture-of-Experts Language Models

### v2

@@ 1 Introduction
Empirical scaling laws for language model pretraining (kaplan_scale; hoffmann2022training; openai2023gpt4; openai2024openai; geminiteam2023gemini; henighan2020scaling; clark2022unified; yun2024toward; ludziejewski2024scaling) have demonstrated that proportionally increasing model capacity, along with data and total compute budget, consistently decreases pretraining loss (i.e., perplexity), improves downstream task performance (devlin_acl_2019_bert; brown_neurips2020_few_shot; srivastava2023beyond) and unlocks emergent capabilities (wei2022emergent).
A recurring notion in these studies is that model capacity is well quantified by the total number of model parameters. However, the number of parameters is not the only means to increase model capacity— compute per example (i.e., a fixed-sized input), measured in FLoating OPerations (FLOPs), also plays a significant role (clark2022unified). In fact, several mechanisms (shazeer2017; dehghani2018universal; wei2022chain_cot; goyal2024think_pause; Csordas2024MoEUTMU) allow for independent variation of the number of parameters or FLOPs per example within a model. For instance, Sparse Mixture-of-Experts (MoE) models (shazeer2017) introduce “FLOP-free parameters” by leveraging sparsity, where only a subset of expert modules is activated for each input.
When studying scaling laws for specific classes of models, e.g., vanilla transformers, the total number of parameters can serve as a reasonable relative estimator of FLOPs per example. Therefore, using the number of parameters as a measure of model capacity in scaling law studies is appropriate. However, in scenarios or for architectures where the number of parameters and FLOPs per example are not directly linked, it is essential to jointly consider the effects of these variables on scaling model capacity (clark2022unified). We therefore ask
@@
To address this question, we study sparse Mixture-of-Expert Transformers (MoEs) (shazeer2017; lepikhin2021gshard; fedus_switch; zoph2022st; muennighoff2024olmoe) in the context of language modeling. Existing scaling law studies for MoEs, investigate the role of variables like number and granularity (ludziejewski2024scaling)of experts, underlying dense model size and inference compute in predicting the performance of the models under different conditions such as training or inference compute optimality (Du2021GLaMES; clark2022unified; yun2024toward; ludziejewski2024scaling). In this paper, we focus on the interaction between FLOPs per example and total parameter count, and their impact on model performance in MoEs, through a large-scale empirical study.
@@
  • During inference, FLOPs per example seem to play a more important role [^1] A relevant discussion here is the recent trend of increasing test-time compute, e.g., OpenAI o1 model (openai2024openai), achieved by generating more tokens as a way for introducing parameter-free-FLOPs.. For many tasks, upstream performance is a good predictor of downstream performance and the relationship between upstream and downstream performance is not impacted by the sparsity level. However, we observe that for models with the same perplexity on the pretraining data distribution, sparser models, i.e., models with fewer number of active parameters, perform worse on specific types of downstream tasks that presumably require more “reasoning”.
Our results, in line with findings from previous relevant studies (ludziejewski2024scaling; he2024mixture) on scaling laws for MoEs, show increasing sparsity level leads to better performance and efficiency during pretraining. Considering the various methods to increase compute per example during inference adaptively conditioned on task or example complexity, we conclude that MoEs approaches, which reduce the unit compute cost (i.e., FLOPs per token) by increasing the sparsity level, hold significant promise given their potential to enhance efficiency in both pretraining and inference.
(a) IsoFLOP surface over sparsity and total parameters
(b) IsoFLOP surface over sparsity and active parameters
Figure 1: IsoFLOP surface over observed pretraining loss $\mathbf{L}$, model size (in terms of total $\mathbf{N}$ and active parameters $\mathbf{N_{a}}$), and sparsity $\mathbf{S}$. We fit a polynomial function mapping $\mathbf{N}$ (or $\mathbf{N_{a}}$), $\mathbf{S}$, and their interaction to $\mathbf{L}$, using empirical data. Both fits achieve an MSE loss of 0.0001 on a held-out set. These results indicate that for a fixed compute budget, increasing model sparsity leads to a reduction in pretraining loss. When considering optimal model size, we observe opposite trends for total parameters ($\mathbf{N}$) (Figure a) versus active parameters ($\mathbf{N_{a}}$) (Figure b). (Figure 8 includes results Section D.1 for different total compute budgets $C$.)
@@ 2 The Interplay between Model Parameters and Sparsity in MoEs
Intuitively, under infinite data setting, scaling model capacity along with the training compute budget leads to performance improvements. Previous scaling law studies suggest that, conditioned on a training compute budget measured in FLOPs denoted by $C$, the optimal number of parameters, $N^{*}(C)$, exhibits a power-law relationship with $C$ (hoffmann2022training):
@@
As the first step, considering a fixed training compute budget $C$, we fit a 3D surface, referred to as the IsoFLOP surface, in Figure 1a, using a polynomial function, following approach II of hoffmann2022training. Compared to hoffmann2022training we include the sparsity variable and fit a single 3d IsoFLOP surface across all data points, rather than fitting separate 2d IsoFLOP curves for fixed sparsity levels or model sizes. We conducted a grid search to determine the optimal polynomial degree for $N$, $S$, and the interaction term $N\times S$, finding that a degree of $(2,2,2)$ resulted in the lowest cross-validation error. Both $N$ and $S$ are in log space (see Appendix B for more details).
As seen in Figure 1a, the IsoFLOP surface plot is parabolic along model size, suggesting that the findings of hoffmann2022training extend to MoEs across different sparsity levels, i.e., $\mathcal{L}(N;C,S)$ is parabolic, with its optimal solution located at the turning point. When considering the total number of parameters $N$, the optimal value increases as the sparsity level increases, while for the active number of parameters $N_{a}$ the optimal value decreases with the sparsity level. This indicates that by increasing the sparsity level the training compute optimal models are larger but have fewer FLOPs per example, i.e., lower inference cost. Moreover, along sparsity, the pretraining loss decreases monotonically, indicating that, for the same compute budget, sparser models achieve better pretraining performance. We observe the same pattern across different training compute budgets (See Section D.1). To better understand and explain these observations, we examine slices of the IsoFLOP surface along the axes of $S$ and $N$ separately in §2.1 and §2.2, respectively.
@@ 3 Impact of Training Compute Budget on the Interaction between Model Parameters and Sparsity
In Figure 3a and Figure 3b, , we see a consistent trend of increasing $N$ and decreasing $N_{a}$ for compute optimal models as sparsity level increases across all training compute budgets. Moreover, as can be seen in Figure 4, when model size in terms of total number of parameters is fixed, optimal sparsity level increases with training compute budget as well as model size as discussed in Section 2.2.
@@ 4 Effect of MoE Sparsity on Downstream Task Performance
If fewer FLOPs per example are the reason behind the worse transfer performance in sparser models, this effect might diminish with a larger total training compute budget, as the optimal active number of parameters increases. Moreover, one can use approaches like chain-of-thought reasoning (wei2022chain_cot) to independently increase FLOPs per example during inference time.
@@ 5 Incorporating Sparsity into Scaling Laws
The scaling laws proposed by kaplan_scale provide a framework for predicting loss in dense models by establishing a power-law relationship between loss $L$, number of parameters $N$ and dataset size $D$, where $N$ and $D$ interact linearly. Formally, the relationship is given by:
@@
The term $\left(1-S\right)$ in the above equation provides a rough estimate of the percentage of number of active parameters. If the exponent for the multiplicative terms is the same then that term provides an approximate estimate of the number of active parameters.
By incorporating sparsity into the scaling law equation, we can eliminate the need for parameters specific to MoEs, such as the total and active number of experts. As demonstrated by frantar2024scaling, this formulation also holds for other sparsity mechanisms, such as weight sparsity, where individual neural network connections are pruned.
We use the recipe described by hoffmann2022training and use the L-BFGS algorithm to fit the coefficients in equation 6 using a Huber loss with $\delta=10^{-3}$. Optimal coefficient values were determined through a grid search (see Table 2 for search values). The results of data fitting and validation are shown in Figure 6. The estimated values are shown in Table 3 in Appendix E.
@@
Figure 6: Scaling law fit on data obtained from training compute-optimal models. Figure 6(a) shows the fit on the data used to estimate the coefficients for equation 6, while Figure 6(b) validates these coefficients on a held-out dataset. All data points with $S=0.98$ were excluded from the fitting process for out-of-sample validation.
@@ 6 Discussion
Our findings amplify the findings of ludziejewski2024scaling and further justify the effort to work toward MoEs with experts larger in number and smaller in size (he2024mixture). For downstream tasks which their performance is predictable given the pretraining loss (i.e., perplexity), sparsity potentially provides efficiency gains both during pretraining and inference.
@@
  • Effect of Sparsity on Downstream Performance: For most downstream tasks, models with similar pretraining perplexity have similar downstream task performance regardless of sparsity. For reading comprehension tasks (e.g., CoQA (reddy-etal-2019-coqa), SQuAD (rajpurkar-etal-2018-know)), denser models perform better, potentially due to their higher inference-time compute than a perplexity-matched sparse model. Strategies to increase inference time compute dynamically (wei2022chain_cot; goyal2024think_pause) may address this gap.
  • Parametric Scaling Law: We propose a parametric form for scaling laws that accounts for sparsity. The model coefficients are estimated using the empirical data obtained by training compute-optimal models. An interesting observation from  Appendix E is that the exponent for sparsity term $\lambda$ is negative which is consistent with our intuition that sparser models lead to a lower perplexity.
@@ 6.1 Limitations
In our analysis, similar to other scaling law studies (kaplan_scale; hoffmann2022training), we have measured the costs for both training and inference exclusively in terms of FLOPs. While there may be discrepancies between actual computational costs and theoretical FLOPs due to hardware specifications, infrastructure, and implementation details, it is reasonable to abstract away from these factors when comparing similar models under fixed conditions. However, an important aspect not accounted for in this study is the cost associated with memory usage and communication overhead, which could potentially increase as we raise the sparsity level. Incorporating these factors is challenging because they are highly dependent on the hardware used. To address this limitation to some extent, in Section 2.2 we investigate the optimal sparsity level under the setting where total number of parameters is fixed.
Despite the limitation with using an approximate method to quantify FLOPs, our findings highlight the importance of investing in methods to enhance the efficiency of sparse Mixture-of-Experts models. By increasing model capacity through additional parameters while minimizing per-unit computation costs, these models have the potential to improve both efficiency and performance. The availability of GPUs with larger memory, for e.g., the recently introduced H200 GPU chip with 141 GB of memory as well as improving the efficiency of training and deployment pipelines (nv_nemo) suggest that there is significant interest in developing efficient implementations for MoEs.
@@ 7.1 Scaling Laws for Language Models
Scaling laws have proven to be a powerful framework for understanding and predicting the performance of language models. Existing studies, such as kaplan_scale and hoffmann2022training, reveal that power-law relationships govern model performance as a function of factors like model size, data size, and compute budget, offering predictable performance improvements with increased resources.
hoffmann2022training emphasizes the critical balance between model size and the number of training tokens when the training compute budget is fixed, showing that scaling the model without corresponding data increases can lead to suboptimal performance. Additionally, Bi2024DeepSeekLS explores more nuanced scaling behaviors by incorporating data quality, demonstrating that higher-quality data allows for more efficient scaling, and thus, a larger portion of the compute budget should be allocated to increasing model size.
Recent work extends scaling law analysis to specialized contexts, including over-training (Gadre2024overtrain), downstream task performance, and multilingual or multi-modal settings, where scaling laws provide valuable insights and can be adapted to address specific challenges.
@@ 7.2 Scaling Laws for MoEs
Mixture-of-Experts (MoE) models (shazeer2017; lepikhin2021gshard; fedus_switch; deepseekr1) have emerged as a powerful architecture for language modeling, primarily because they decouple computational cost from parameter count. This separation between parameters and FLOPs per token in MoE architectures calls for scaling laws that can accurately factor in the contributions of both.
Previous research on the scaling behavior of MoE models has established foundational scaling laws, incorporating factors such as total parameter count, the number of experts, and the granularity of these experts (clark2022unified; ludziejewski2024scaling; wang-etal-2024-scaling). However, these studies typically assume a fixed configuration for other critical variables influencing FLOPs per token, such as the number of active experts per input. In contrast, we propose a generalized scaling law that considers variables like active parameter count and sparsity level, thereby expanding the applicability of MoE scaling laws.
@@
Supporting this, Du2021GLaMES demonstrates GLaM’s superior efficiency and performance compared to GPT-3, showing that MoE architectures can achieve high performance with significantly lower computational and energy costs. Further insights are offered by clark2022unified, who analyze scaling behaviors across various MoE routing techniques. While their study finds that MoEs generally outperform dense models, it also notes diminishing benefits as base model sizes grow. ludziejewski2024scaling challenge this conclusion, attributing the diminished returns partly to the fixed number of training tokens across models and constant expert sizes. By introducing "granularity" and adjusting training durations, they demonstrate that MoEs can outperform dense models across any compute budget, debunking the notion of diminishing returns for MoEs with adaptive expert configurations. More recently, jelassi2024mixture finds that, on downstream tasks, MoEs scale efficiently with the number of experts (i.e., increasing sparsity) on memorization tasks, but their reasoning capabilities saturate and lag behind dense models on tasks requiring complex reasoning when compared based on total number of parameters.
Another approach by he2024mixture explores the benefits of training MoEs with larger numbers of smaller experts rather than the conventional setup of fewer, larger experts. They introduce Parameter Efficient Expert Retrieval (PEER), a novel routing mechanism designed to tackle the computational and optimization challenges that arise when handling a high number of experts, thus enabling efficient scaling of MoE models.
Lastly, yun2024toward draws attention to the increased inference costs associated with scaling MoEs by adding experts. While additional experts may not substantially affect training costs, they can inflate inference costs, thereby diminishing deployment efficiency. To address this, the study proposes an over-trained budget allocation strategy, optimizing MoE models for both performance and efficiency in deployment.

### v3

# Understanding Compute-Parameter Trade-offs in Sparse Mixture-of-Experts Language Models

@@ 1 Introduction
Empirical scaling laws for language model pretraining (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2023, 2024; Gemini Team et al., 2024; Henighan et al., 2020; Clark et al., 2022; Yun et al., 2024; Ludziejewski et al., 2024) have demonstrated that proportionally increasing model capacity, along with data and total compute budget, consistently decreases pretraining loss (i.e., perplexity), improves downstream task performance (Devlin et al., 2019; Brown et al., 2020; BIG-bench authors, 2023) and unlocks emergent capabilities (Wei et al., 2022a).
A recurring notion in these studies is that model capacity is well quantified by the total number of model parameters. However, the number of parameters is not the only means to increase model capacity. Compute per example (i.e., a fixed-sized input), measured in FLoating OPerations (FLOPs), also plays a significant role (Clark et al., 2022). In fact, several mechanisms (Shazeer et al., 2017; Dehghani et al., 2019; Wei et al., 2022b; Goyal et al., 2024; Csord’as et al., 2024) allow for independent variation of the number of parameters or FLOPs per example within a model. For instance, Sparse Mixture-of-Experts (MoE) models (Shazeer et al., 2017) introduce “FLOP-free parameters” by leveraging sparsity, where only a subset of expert modules is activated for each input.
(a) IsoFLOP surface over sparsity and total parameters
(b) IsoFLOP surface over sparsity and active parameters
Figure 1: IsoFLOP surface over observed pretraining loss $\mathbf{L}$, model size (in terms of total $\mathbf{N}$ and active parameters $\mathbf{N_{a}}$), and sparsity $\mathbf{S}$. We fit a polynomial function mapping $\mathbf{N}$ (or $\mathbf{N_{a}}$), $\mathbf{S}$, and their interaction to $\mathbf{L}$, using empirical data. For both fits the MSE loss for predicting loss on a held out set is $0.0001$. These results indicate that for a fixed compute budget, increasing model sparsity leads to a reduction in pretraining loss. When considering optimal model size, we observe opposite trends for total parameters ($\mathbf{N}$) (Figure a) versus active parameters ($\mathbf{N_{a}}$) (Figure b). (See Figure 8 in Section D.1 for results with different total compute budgets $C$.)
When studying scaling laws for specific classes of models, e.g., vanilla transformers, the total number of parameters can serve as a reasonable relative estimator of FLOPs per example. Therefore, using the number of parameters as a measure of model capacity in scaling law studies is appropriate. In scenarios or for architectures where the number of parameters and FLOPs per example are not directly linked, it is essential to jointly consider the effects of these variables on scaling model capacity (Clark et al., 2022). We therefore ask
@@
To address this question, we study sparse Mixture-of-Expert Transformers (MoEs) (Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022; Zoph et al., 2022; Muennighoff et al., 2024) in the context of language modeling. Existing scaling law studies for MoEs, investigate the role of variables like number and granularity (Ludziejewski et al., 2024) of experts, underlying dense model size and inference compute in predicting the performance of the models under different conditions such as training or inference compute optimality (Du et al., 2021; Clark et al., 2022; Yun et al., 2024; Ludziejewski et al., 2024). In this paper, we focus on the interaction between FLOPs per example and total parameter count, and their impact on model performance in MoEs, through a large-scale empirical study.
@@
  • During inference, FLOPs per example seem to play a more important role [^1] A relevant discussion here is the recent trend of increasing test-time compute, e.g., OpenAI o1 model (OpenAI, 2024), achieved by generating more tokens as a way for introducing parameter-free-FLOPs.. For many tasks, upstream performance is a good predictor of downstream performance and the relationship between upstream and downstream performance is not impacted by the sparsity level. However, on downstream tasks that presumably require more “reasoning”, we observe that for models with the same perplexity on the pretraining data distribution, sparser models, i.e., models with fewer active parameters, perform worse.
Our results, in line with findings from previous relevant studies (Ludziejewski et al., 2024; He, 2024) on scaling laws for MoEs, show increasing sparsity level leads to better performance and efficiency during pretraining. Considering the various methods to increase compute per example during inference adaptively conditioned on task or example complexity, we conclude that approaches like MoEs, which reduce the unit compute cost (i.e., FLOPs per token) by increasing the sparsity level, hold significant promise given their potential to enhance efficiency in both pretraining and inference.
@@ 2 The Interplay between Model Parameters and Sparsity in MoEs
Intuitively, under infinite data setting, scaling model capacity along with the training compute budget leads to performance improvements. Previous scaling law studies suggest that, conditioned on a training compute budget measured in FLOPs denoted by $C$, the optimal number of parameters, $N^{*}(C)$, exhibits a power-law relationship with $C$ (Hoffmann et al., 2022):
@@
As the first step, considering a fixed training compute budget $C$, we fit a 3D surface, referred to as the IsoFLOP surface, in Figure 1a, using a polynomial function, following approach II of Hoffmann et al. (2022). Compared to Hoffmann et al. (2022) we include the sparsity variable and fit a single 3d IsoFLOP surface across all data points, rather than fitting separate 2d IsoFLOP curves for fixed sparsity levels or model sizes. We conducted a grid search to determine the optimal polynomial degree for $N$, $S$, and the interaction term $N\times S$, finding that a degree of $(2,2,2)$ resulted in the lowest cross-validation error. Both $N$ and $S$ are in log space (see Appendix B for more details).
As seen in Figure 1a, the IsoFLOP surface plot is parabolic along model size, suggesting that the findings of Hoffmann et al. (2022) extend to MoEs across different sparsity levels, i.e., $\mathcal{L}(N;C,S)$ is parabolic, with its optimal solution located at the turning point. When considering the total number of parameters $N$, the optimal value increases as the sparsity level increases, while for the active number of parameters $N_{a}$ the optimal value decreases with the sparsity level. This indicates that by increasing the sparsity level the training compute optimal models are larger but have fewer FLOPs per example, i.e., lower inference cost. Moreover, along sparsity, the pretraining loss decreases monotonically, indicating that, for the same compute budget, sparser models achieve better pretraining performance. We observe the same pattern across different training compute budgets (See Section D.1). To better understand and explain these observations, we examine slices of the IsoFLOP surface along the axes of $S$ and $N$ separately in §2.1 and §2.2, respectively.
@@ 3 Impact of Training Compute Budget on the Interaction between Model Parameters and Sparsity
In Figure 3a and Figure 3b, , we see a consistent trend of increasing $N$ and decreasing $N_{a}$ for compute optimal models as sparsity level increases across all training compute budgets. Moreover, as can be seen in Figure 4, when model size in terms of total number of parameters is fixed, optimal sparsity level decreases with training compute budget while increases with model size as discussed in Section 2.2.
@@ 4 Effect of MoE Sparsity on Downstream Task Performance
If fewer FLOPs per example are the reason behind the worse transfer performance in sparser models, this effect might diminish with a larger total training compute budget, as the optimal active number of parameters increases. Moreover, one can use approaches like chain-of-thought reasoning (Wei et al., 2022b) to independently increase FLOPs per example during inference time.
In Appendix E, we explore whether increasing inference-time compute via Chain-of-Thought (CoT) prompting can improve the performance of MoEs on tasks that require more reasoning. Our experiments indicate that MoEs benefit more from this increased compute compared to dense models with a similar number of active parameters. This suggests that dynamic compute allocation during inference may be crucial for MoEs to perform well on complex reasoning tasks.
@@ 5 Incorporating Sparsity into Scaling Laws
The scaling laws proposed by Kaplan et al. (2020) provide a framework for predicting loss in dense models by establishing a power-law relationship between loss $L$, number of parameters $N$ and dataset size $D$, where $N$ and $D$ interact linearly. Formally, the relationship is given by:
@@
The term $\left(1-S\right)$ in the above equation provides a rough estimate of the percentage of active parameters. If the exponent for the multiplicative terms is the same then that term provides an approximate estimate of the number of active parameters.
By incorporating sparsity into the scaling law equation, we can eliminate the need for parameters specific to MoEs, such as the total and active number of experts. As demonstrated by Frantar et al. (2024), this formulation also holds for other sparsity mechanisms, such as weight sparsity, where individual neural network connections are pruned.
We use the recipe described by Hoffmann et al. (2022) and use the L-BFGS algorithm to fit the coefficients in equation 6 using a Huber loss with $\delta=10^{-3}$. Optimal coefficient values were determined through a grid search (see Table 2 for search values). The results of data fitting and validation are shown in Figure 6. The estimated values are shown in Table 3 in Appendix F.
@@
Figure 6: Scaling law fit on data obtained from training compute-optimal models. Figure 6(a) shows the fit on the data used to estimate the coefficients for equation 6, while Figure 6(b) validates these coefficients on a held-out dataset. All data points with $S=0.98$ were excluded from the fitting process for out-of-sample validation. The dashed lines represent equal loss values.
@@ 6 Discussion
Our findings amplify the findings of Ludziejewski et al. (2024) and further justify the effort to work toward MoEs with experts larger in number and smaller in size (He, 2024). For downstream tasks which their performance is predictable given the pretraining loss (i.e., perplexity), sparsity potentially provides efficiency gains both during pretraining and inference.
@@
  • Effect of Sparsity on Downstream Performance: For most downstream tasks, models with similar pretraining perplexity have similar downstream task performance regardless of sparsity. For reading comprehension tasks (e.g., CoQA (Reddy et al., 2019), SQuAD (Rajpurkar et al., 2018)), denser models perform better, potentially due to their higher inference-time compute than a perplexity-matched sparse model. Strategies to increase inference time compute dynamically (Wei et al., 2022b; Goyal et al., 2024) may address this gap.
  • Parametric Scaling Law: We propose a parametric form for scaling laws that accounts for sparsity. The model coefficients are estimated using the empirical data obtained by training compute-optimal models. An interesting observation from  Appendix F is that the exponent for sparsity term $\lambda$ is negative which is consistent with our intuition that sparser models lead to a lower perplexity.
@@ 6.1 Limitations
In our analysis, similar to other scaling law studies (Kaplan et al., 2020; Hoffmann et al., 2022), we have measured the costs for both training and inference exclusively in terms of FLOPs. While there may be discrepancies between actual computational costs and theoretical FLOPs due to hardware specifications, infrastructure, and implementation details, it is reasonable to abstract away from these factors when comparing similar models under fixed conditions. However, an important aspect not accounted for in this study is the cost associated with memory usage and communication overhead, which could potentially increase as we raise the sparsity level. Incorporating these factors is challenging because they are highly dependent on the hardware used. To address this limitation to some extent, in Section 2.2 we investigate the optimal sparsity level under the setting where total number of parameters is fixed.
Despite the limitation with using an approximate method to quantify FLOPs, our findings highlight the importance of investing in methods to enhance the efficiency of sparse Mixture-of-Experts models. By increasing model capacity through additional parameters while minimizing per-unit computation costs, these models have the potential to improve both efficiency and performance. The availability of GPUs with larger memory, for e.g., the recently introduced H200 GPU chip with 141 GB of memory as well as improving the efficiency of training and deployment pipelines (NeMo Authors, 2025) suggest that there is significant interest in developing efficient implementations for MoEs.
@@ 7.1 Scaling Laws for Language Models
Scaling laws have proven to be a powerful framework for understanding and predicting the performance of language models. Existing studies, such as Kaplan et al. (2020) and Hoffmann et al. (2022), reveal that power-law relationships govern model performance as a function of factors like model size, data size, and compute budget, offering predictable performance improvements with increased resources.
Hoffmann et al. (2022) emphasizes the critical balance between model size and the number of training tokens when the training compute budget is fixed, showing that scaling the model without corresponding data increases can lead to suboptimal performance. Additionally, DeepSeek-AI (2024) explores more nuanced scaling behaviors by incorporating data quality, demonstrating that higher-quality data allows for more efficient scaling, and thus, a larger portion of the compute budget should be allocated to increasing model size.
Recent work extends scaling law analysis to specialized contexts, including over-training (Gadre et al., 2024), downstream task performance, and multilingual or multi-modal settings, where scaling laws provide valuable insights and can be adapted to address specific challenges.
@@ 7.2 Scaling Laws for MoEs
Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022; DeepSeek-AI, 2025) have emerged as a powerful architecture for language modeling, primarily because they decouple computational cost from parameter count. This separation between parameters and FLOPs per token in MoE architectures calls for scaling laws that can accurately factor in the contributions of both.
Previous research on the scaling behavior of MoE models has established foundational scaling laws, incorporating factors such as total parameter count, the number of experts, and the granularity of these experts (Clark et al., 2022; Ludziejewski et al., 2024; Wang et al., 2024). However, these studies typically assume a fixed configuration for other critical variables influencing FLOPs per token, such as the number of active experts per input. In contrast, we propose a generalized scaling law that considers variables like active parameter count and sparsity level, thereby expanding the applicability of MoE scaling laws.
@@
Supporting this, Du et al. (2021) demonstrates GLaM’s superior efficiency and performance compared to GPT-3, showing that MoE architectures can achieve high performance with significantly lower computational and energy costs. Further insights are offered by Clark et al. (2022), who analyze scaling behaviors across various MoE routing techniques. While their study finds that MoEs generally outperform dense models, it also notes diminishing benefits as base model sizes grow. Ludziejewski et al. (2024) challenge this conclusion, attributing the diminished returns partly to the fixed number of training tokens across models and constant expert sizes. By introducing "granularity" and adjusting training durations, they demonstrate that MoEs can outperform dense models across any compute budget, debunking the notion of diminishing returns for MoEs with adaptive expert configurations. More recently, Jelassi et al. (2024) finds that, on downstream tasks, MoEs scale efficiently with the number of experts (i.e., increasing sparsity) on memorization tasks, but their reasoning capabilities saturate and lag behind dense models on tasks requiring complex reasoning when compared based on total number of parameters.
Another approach by He (2024) explores the benefits of training MoEs with larger numbers of smaller experts rather than the conventional setup of fewer, larger experts. They introduce Parameter Efficient Expert Retrieval (PEER), a novel routing mechanism designed to tackle the computational and optimization challenges that arise when handling a high number of experts, thus enabling efficient scaling of MoE models.
Lastly, Yun et al. (2024) draws attention to the increased inference costs associated with scaling MoEs by adding experts. While additional experts may not substantially affect training costs, they can inflate inference costs, thereby diminishing deployment efficiency. To address this, the study proposes an over-trained budget allocation strategy, optimizing MoE models for both performance and efficiency in deployment.